A fir tree is 5 times as old as a pine tree. Seven years from now , the sum of their ages will be 32. How old is the fir tree now?
Answer by darq(90)

f=5p
(f+7)+(p+7)=32

substitute f with 5p in the second equation because according to the first equation they are same.

5p+7+p+7=32
6p+14=32
6p=18
pine=3 years old

fir is five times older so it is 3*5=15 years old.